OpenStudy (anonymous):

If g(x) = x2 − 3, find g(4). 13 5 16 −3

OpenStudy (zale101):

g(x)=x^2-3 g(4)=4^2-3 = ? g of 4 means that you're replacing 4 by the variable

OpenStudy (skullpatrol):

Actually, you replace the variable with the numeral 4.

OpenStudy (anonymous):

is it 4-3

OpenStudy (anonymous):

or 4^2-3

OpenStudy (skullpatrol):

16 - 3

OpenStudy (skullpatrol):